(7) An employee joined a company in 2017 with a starting salary of 50000 USD. Every year this employee receives a raise of 1000 USD plus 5 percent of the salary of the previous year. Let Cn denote the employee's salary n years after 2017.
(a) Set up a recurrence relation for Cn.
(b) What will the salary of this employee be in 2025?
(c) Find an explicit formula for Cn.
1
Expert's answer
2020-10-17T16:42:04-0400

C0=50 000

a.)Cn=Cn-1*1.05 + 1 000

b.)C1=53 500

C2=57 175

C3=61 033.75

C4=65 085.4375

C5=69 339.7094

C6=73 806.6949

C7=78 497.0296

C2025-2017=C8=83 421.8811  

c.) "C=50000"

"k= 1 + 5\/100"

"d=1000"

"C"1"=C*k+d,"

"C"2"=C"1"*k+d=C*k^2+d*k+d" "="

"C*k^2+d*(k+1),"

"C"3"=C*k^3+d*(k^2+k+1),"

"C_4" "=C*k^4+d*(k^3+k^2+k+1),"

"\sum_{i=0}^{n-1} k^i=(k^n-1)/(k-1),"

Cn"=C*k^n+d*(k^n-1)\/(k-1)=50000*1.05^n+1000*(1.05^n-1)\/0.05"